# Tracing the Techno-Supremacy Doctrine: A Critical Discourse Analysis of the AI Executive Elite

### v1

@@ 4 The Spectrum of TSD Discourse
Figure 2 shows the spectrum of stances in terms of TSD Adherence and Balance. On the Y-axis (TSDA), higher positive values mean more pro-TSD leaning discourse, whereas negative values mean more anti-TSD leaning discourse. The X-axis (TSDB) ranges from 0, corresponding to completely one-sided discourse (either pro- or anti-TSD), to 0.5, representing perfectly balanced discourse. The top-left region (high TSDA, low TSDB) represents imbalanced, strongly pro-TSD texts. The bottom-left region (negative TSDA, low TSDB) represents imbalanced, strongly anti-TSD texts. [^4] Note that the region along the Y-axis where TSDB is high will necessarily contain texts with TSDA scores close to zero. This is because a TSDB value close to 0.5 means that the values of the pro- and anti-TSD components of TSDA are very similar, resulting in their near or total cancellation. Therefore, it is impossible to find texts in either the top-right or bottom-right regions, as highly balanced texts will inherently have low TSDA and cluster around the X-axis.
The typical text in the corpus tends to be moderately pro-TSD, presenting this stance in a relatively imbalanced way (TSDA average: 13.37, TSDB average: 0.24). Notably, however, both TSDA and TSDB have very high standard deviations (22.49 and 0.17, respectively) considering their ranges (approximately [-23, 55] and [0, 0.5], respectively). Therefore, the AI executive elite is not a monolithic group when it comes to how TSD is expressed and how AI-related concerns are responded to. While, on average, texts tend to be pro-TSD, some texts are strongly pro-TSD (e.g., altman-2024-intelligence), strongly anti-TSD (e.g., amodei-2023-scaling), or much more neutral (e.g., khosla-2017-scary). Similarly, while pro- or anti-TSD discourse is typically moderately more dominant within a given text, some texts are notably imbalanced (e.g., andreessen-2023-manifesto), while others are remarkably balanced (e.g., taneja-2022-consequences).
Texts cluster into three main profiles. The first profile (top-left to mid-left on the chart) is characterized by high positive TSDA scores, indicating a strong pro-TSD stance, and relatively low TSDB scores, meaning that this pro-TSD stance is presented in a fairly imbalanced way. Both of Andreessen’s texts are extreme examples of this one-sided pro-TSD discourse (TSDB near 0). The second profile (bottom half of the chart) contains critical and counter-TSD texts. amodei-2023-scaling is an outlier, with a very strongly negative TSDA score and a low TSDB score, indicating that it is not only heavily anti-TSD in its overall message but also quite imbalanced in that direction. suleyman-2024-containment also has a negative TSDA but a moderate TSDB, indicating an anti-TSD leaning achieved with more internal balance, but still weighted towards critique. Both pichai-2020-regulate and khosla-2017-scary are near the neutral TSDA axis with high TSDB scores, which suggests their slight anti-TSD leaning is achieved through a relatively balanced presentation of pro- and anti-TSD elements. The third profile (right side of the chart) contains balanced texts, where both TSDA components wholly or nearly cancel each other out (e.g., khosla-2017-scary, taneja-2022-consequences).
@@
Discourse reproducing TSD’s core tenets (cf. TCE codes) is more voluminous than TSD-reinforcing responses to AI-related concerns (cf. TRR codes). Conversely, anti-TSD responses to AI-related concerns (cf. ANTI-TRR codes) are more prevalent than explicit contradictions of core TSD aspects (cf. ANTI-TCE codes), which are extremely rare. Two specific TCE aspects are by far the most common across the corpus: believing that technology will create a better, even utopian world (CT-UF) and framing technology as the primary or inevitable engine driving positive societal transformation (TC-PD). Conversely, the most prevalent ANTI-TRR aspect across the corpus is explicitly acknowledging a concern (ACK-RI); however, as will be discussed later, this is usually used as a mere preamble to pro-TSD discourse.
@@
The analysis of TSDA by sub-component reveals a distinct TSD manifestation, referred to here as Benign Techno-Optimism (BTO). This type of discourse is characterized by a high degree of Cult of Technology expressions, particularly regarding utopian futures (CT-UF) and the almost magical power of technology (CT-MP). However, these optimistic expressions are critically tempered by significant counterarguments to TSD. These arguments include direct contradictions to TSD tenets (cf. ANTI-TCE codes), especially those challenging the idea of technology as an inevitable driver of progress (ANTI-TC-PD), as well as critical responses to AI-related concerns (cf. ANTI-TRR codes), particularly the proposal of non-technological solutions (ADD-SN). suleyman-2024-containment and amodei-2024-grace exemplify BTO, displaying an overall optimism and enthusiasm about AI’s potential, critically tempered with a combination of direct contradictions of core TSD tenets and discussion of non-technological approaches.
@@
The remainder of the section presents a discussion on two discursive phenomena. These phenomena were selected for their high frequency in the quantitative analysis and their central importance to the reproduction of TSD.
4.1 AI-Driven Utopia
@@
The analysis uncovered that believing in an (almost) utopian future (CT-UF) and considering technology the main or inevitable driver of societal progress (TC-PD) are not only the most expressed core TSD aspects but also co-occur in ten of the 14 analyzed texts. This discursive strategy creates a compelling narrative: It first portrays AI as the inevitable culmination of technological progress (TC-PD), and then heralds an era of unparalleled, universally shared prosperity driven by that progress (CT-UF). This dual assertion of historical inevitability plus utopian promise functions to legitimize current AI development and deployment paradigms, justifying immense investment and focus (ADD-JU). Additionally, this narrative can be used to marginalize critical perspectives or calls for caution as “anti-progress” or unduly pessimistic (cf. “AI doomers”). In doing so, it subtly downplays the intricate requirements of social adaptation, robust governance, and inclusive human agency essential for the responsible development of AI.
4.2 Pro-TSD Risk Acknowledgment
@@
While direct acknowledgment of risk (ACK-RI) is the most common anti-TSD move, the data reveals that it was never the final word in any of the analyzed texts. In 13 of the 14 texts, ACK-RI discourse was followed by a discursive move aimed at managing, justifying, or resolving the concern in a pro-TSD manner (cf. TRR codes). The single text where this did not occur is andreessen-2023-manifesto, which contained no ACK-RI discourse. The combination of ACK-RI with ADD-ST is particularly insidious, as it leads to a subtle reinforcement of TSD. The initial acknowledgment validates the reader’s concerns, but the immediate pivot to technological fixes implicitly frames these risks as ultimately problems that can be solved by further technological innovation. The insidious nature of this strategy lies in its power of naturalization, a key mechanism of ideological power in Fairclough’s framework. By framing the conversation as ‘risk acknowledged, tech solution proposed,’ the discourse makes the pivot to a technological fix seem common sense and natural. When repeated over time, this pattern works to naturalize the underlying assumption that the appropriate response to technological harm is always more or better technology, effectively marginalizing non-technological solutions until they seem impractical or irrelevant. The ideology becomes most potent when it is no longer perceived as ideology, but simply as the way things are naturally done. The presence of ACK-RI poses a challenge when tracing TSD in discourse. While ACK-RI might initially suggest a balanced and critical perspective, its communicative function can be multifaceted and is often used as a preamble to pro-TSD discourse. A reader encountering ACK-RI discourse may unconsciously adopt a position of engagement with the text, making subsequent pro-TSD arguments more palatable. Distinguishing genuine, critical engagement from pro-TSD rhetorical maneuvering is vital. Truly anti-TSD discourse (e.g., BTO) would use ACK-RI as a preamble to comprehensively explore diverse solutions, including non-tech proposals (ADD-SN), and/or question core TSD tenets (cf. ANTI-TCE codes). The actual impact of ACK-RI thus depends on the subsequent discursive moves and the overall balance of solution types proposed.
5 The Dynamics of TSD Discourse
@@
Figure 3 shows TSD Adherence over time. The chart is contextualized with key milestones in technological capability (e.g., the launch of ChatGPT), foundational ethical and policy framings (e.g., the EU AI Act), and international governance dialogues (e.g., the AI Safety Summit). More details about this broader context are provided in the following section, serving as a backdrop for the subsequent analysis.
The data reveals distinct patterns of discursive change over time, with most authors demonstrating significant shifts toward more pronounced and imbalanced pro-TSD stances. For example, Pichai moved from a balanced, slightly anti-TSD position in 2020 (TSDA = -6.35, TSDB = 0.43) to a strongly pro-TSD and imbalanced one in 2025 (TSDA = 28.83, TSDB = 0.18). Similarly, Khosla shifted from a highly balanced, neutral stance in 2017 (TSDA = -0.94, TSDB = 0.49) to a strongly pro-TSD and imbalanced position in 2024 (TSDA = 32.38, TSDB = 0.26). This dynamic can also be rapid. Within the same year, Altman moved from a moderately pro-TSD text (TSDA = 9.51, TSDB = 0.30) to one that is extremely pro-TSD and imbalanced (TSDA = 55.15, TSDB = 0.09).
Beyond these individual trajectories, a broader trend of polarization emerges across the corpus, particularly in the period following the launch of ChatGPT. The period from late 2023 onwards exhibits the most extreme TSDA scores in both positive and negative directions. Both the highest pro-TSD scores observed in the corpus (altman-2024-intelligence, andreessen-2023-manifesto, khosla-2024-utopia, pichai-2025-innovation) and the most significantly anti-TSD scores (amodei-2023-scaling, suleyman-2024-containment) are all concentrated in this recent period. Overall, five of the seven authors (71%) showed an increase in their TSDA scores over time, with four of these five also becoming less balanced. Suleyman is the only author whose TSDA score decreases (and TSDB score increases) after the launch of ChatGPT.
@@
The remainder of the section provides a qualitative analysis of the recent surge of pro-TSD discourse. The broader societal situational context in which relevant texts were produced is described first. Then, a discussion on potential social determinants contributing to this discursive shift is presented.
5.1 From AI Safety to Action
@@
The period following the launch of ChatGPT in late 2022 has witnessed a dramatic acceleration in AI development and a shift in the dominant discourse among the elite. While the initial generative AI boom was characterized by a robust debate on long-term safety and ethics, this was soon challenged and eclipsed by a powerful counter-narrative emphasizing rapid innovation and geopolitical competition. The new American administration epitomized this shift at the global AI Action Summit in early 2025, where Vice President James David Vance explicitly declared that he was there to discuss opportunity, not safety (noa, 2025c).
@@
As AI capabilities surged after the publication of the Transformer architecture in 2017, the demand for ethical guidelines and regulatory frameworks intensified. Building on a decade of several high-level and voluntary ethics principles from corporations (e.g., Google’s (Pichai, 2018)) and organizations (e.g., OECD’s (noa, 2019)), the European Union AI Act, first proposed in 2021, signaled a concrete move towards binding regulation (noa, 2023a). The unprecedented user adoption of ChatGPT put AI into the global public consciousness. Throughout 2023, media coverage increasingly fixated on concerns such as algorithmic bias, the dangers of misinformation, and AI-driven job displacement. These anxieties were amplified by seminal figures like Geoffrey Hinton, who resigned from Google to warn of long-term risks, and by the Future of Life Institute’s open letter, urging to “immediately pause the training of AI systems more powerful than GPT-4” in March 2023 (noa, 2023b). In October 2023, US President Joe Biden issued a comprehensive Executive Order (noa, 2023c), focusing on AI safety ahead of the first global AI Safety Summit hosted by the United Kingdom in November. The summit culminated in the Bletchley Declaration, in which 28 countries, including the US and China, acknowledged the potential for "serious, even catastrophic, harm" and called for international cooperation (noa, 2023d). The subsequent AI Seoul Summit in May 2024 saw 16 leading AI companies make AI safety commitments; however, it subtly foreshadowed the changing consensus by explicitly adding innovation to the global agenda.
@@
A confluence of powerful technological, commercial, and political forces fueled the discursive shift towards action and innovation. Technologically, the past several years have witnessed an explosion of breakthroughs, including the development of increasingly powerful hardware (e.g., Nvidia’s GPU architecture series), models (e.g., OpenAI’s GPT series; Google’s BERT, Bard, LaMDA, and Gemini series; Anthropic’s Claude series), and practical applications (e.g., DeepMind’s AlphaFold series). Commercially, these technological achievements have attracted unprecedented venture capital, with global AI funding surging past \$100 billion in 2024 (Glaser and Yao, 2025). Additionally, there have been high-profile strategic alliances between companies (e.g., Microsoft’s deepened partnership with OpenAI, investments by Google and Amazon in Anthropic) to ensure market dominance in a hyper-competitive environment. Politically, competition with China has instilled a sense of urgency for Western innovation, which has been intensified by the release of high-capability open-source models (e.g., Meta’s Llama series, xAI’s Grok series) and the launch of the R1 model by the Chinese company DeepSeek in early 2025. Domestically within the United States, the election of Donald Trump marked a pivotal moment. His administration swiftly replaced the previous safety-focused Executive Order with a directive aimed at “removing barriers to American leadership in AI” (noa, 2025d). Moreover, it publicly backed an OpenAI-led \$500 billion investment in AI (noa, 2025e). This proactive stance garnered international support at the AI Action Summit, hosted by France in February 2025. This summit solidified a shift in the global consensus, redirecting the primary focus of AI governance discussions from a predominantly safety-based agenda to a broader one that champions action, opportunity, and accelerated innovation.
5.2 The Elite’s Discursive Shift

### v2

@@ 4.1 The Spectrum of TSD Discourse
Figure 2 shows the stances’ spectrum in terms of TSD Adherence and Balance. On the Y-axis (TSDA), higher positive values mean more pro-TSD leaning discourse, whereas negative values mean more anti-TSD leaning discourse. The X-axis (TSDB) ranges from 0, corresponding to completely one-sided discourse (either pro- or anti-TSD), to 0.5, representing perfectly balanced discourse. The top-left region (high TSDA, low TSDB) represents imbalanced, strongly pro-TSD texts. The bottom-left region (negative TSDA, low TSDB) represents imbalanced, strongly anti-TSD texts. [^5] The region along the Y-axis where TSDB is high will necessarily contain texts with TSDA scores close to zero. This is because a TSDB value close to 0.5 indicates that the pro- and anti-TSD components of TSDA are very similar, leading to near or complete cancellation.
The typical text in the corpus tends to be moderately pro-TSD, presenting this stance in a relatively imbalanced way. Notably, however, both TSDA and TSDB have very high standard deviations, indicating that the AI executive elite is not a monolithic group in how TSD is expressed and how AI-related concerns are addressed. While, on average, texts tend to be pro-TSD, some texts are strongly pro-TSD (e.g., altman-2024-intelligence), strongly anti-TSD (e.g., amodei-2023-scaling), or much more neutral (e.g., khosla-2017-scary). Similarly, while pro- or anti-TSD discourse is typically moderately more dominant within a given text, some texts are notably imbalanced (e.g., andreessen-2023-manifesto), while others are remarkably balanced (e.g., taneja-2022-consequences).
Texts cluster into three main profiles. The first profile (top-left to mid-left on the chart) is characterized by a strong pro-TSD stance presented in a fairly imbalanced way. Both of Andreessen’s texts are extreme examples of this one-sided pro-TSD discourse. The second profile (bottom half of the chart) contains critical and counter-TSD texts. amodei-2023-scaling is an outlier, not only heavily anti-TSD in its overall message but also quite imbalanced in that direction. While suleyman-2024-containment also has anti-TSD leaning, it is achieved with more internal balance. Both pichai-2020-regulate and khosla-2017-scary exhibit a slight anti-TSD lean, achieved through a relatively balanced presentation of pro- and anti-TSD elements. The third profile (on the right side of the chart) contains balanced texts (e.g., khosla-2017-scary, taneja-2022-consequences).
@@
Discourse reproducing TSD’s core tenets is more voluminous than TSD-reinforcing responses to AI-related concerns. Conversely, anti-TSD responses to AI-related concerns are more prevalent than explicit contradictions of core TSD aspects, which are extremely rare. Two specific TSD core aspects are by far the most common across the corpus: (i) believing that technology will create a better, even utopian world and (II) framing technology as the primary or inevitable engine driving positive societal transformation. Conversely, the most prevalent response to concerns across the corpus is explicit acknowledgment; however, as will be discussed later, this is often used merely as a preamble to pro-TSD discourse.
@@
The analysis reveals a distinct TSD manifestation, referred to here as Benign Techno-Optimism (BTO). This type of discourse is characterized by a high degree of Cult of Technology expressions, particularly regarding utopian futures (CT-UF) and the almost magical power of technology (CT-MP). However, these optimistic expressions are critically tempered by significant counterarguments to TSD. These arguments include direct contradictions of TSD tenets, especially those challenging the idea of technology as an inevitable driver of progress (ANTI-TC-PD), as well as critical responses to AI-related concerns, particularly proposals for solutions that are not primarily based on technology (ADD-SN). Both suleyman-2024-containment and amodei-2024-grace exemplify BTO.
@@
The remainder of the section presents a discussion on two discursive phenomena selected for their high frequency in the quantitative analysis and their central importance to the reproduction of TSD.
4.1.1 AI-Driven Utopia
Believing in an (almost) utopian future (CT-UF), and considering technology the main or inevitable driver of societal progress (TC-PD) are the most expressed core TSD aspects. Moreover, they are employed together in ten of the 14 analyzed texts (71%).
This discursive strategy is exemplified by altman-2024-intelligence. This text is likely intended to shape public opinion about the trajectory and future of AI. It engages with and builds upon previous narratives about technological progress and economic transformation, aiming to counter dystopian concerns with an overwhelmingly positive vision (e.g., andreessen-2023-manifesto, khosla-2024-utopia). Consider the following excerpt:
With these new [AI] abilities, we can have shared prosperity to an unimaginable degree today; in the future, everyone’s lives can be better than anyone’s life is now.
Here is one narrow way to look at human history: after thousands of years of compounding scientific discovery and technological progress, we have figured out how to melt sand, add some impurities, arrange it with astonishing precision at extraordinarily tiny scale into computer chips, run energy through it, and end up with systems capable of creating increasingly capable artificial intelligence.
This may turn out to be the most consequential fact about all of history so far. (Altman, 2024a, p. 2)
The first sentence presents an AI-driven utopian future, making the reader receptive to the idea that large-scale AI development is desirable. Its declarative nature positions the author as a visionary who understands AI’s true, transformative potential. Agency is implicitly given to AI (“AI abilities”), while humanity (“we”) is portrayed as a passive beneficiary. The sentence highlights betterment (“prosperity,” “better”), its magnitude (“unimaginable,” “everyone’s lives can be better than anyone’s life”), and its universality (“we,” “shared,” “everyone”), preemptively countering criticisms that AI may exacerbate social inequalities. The juxtaposition of “today” and “the future” emphasizes AI’s potential (rather than its shortcomings). The author employs rewording (i.e., expressing the same idea in different ways) by connecting two somewhat redundant and reinforcing claims. Although the author uses modal verbs (“can have,” “can be”) to express potential and capability, the declarative mood of the sentence portrays these possibilities as inevitable or highly likely. The second sentence details a technological development process culminating in AI, which legitimizes it as the logical and natural pinnacle of thousands of years of human endeavor. Agency is explicitly given to humanity (“we have figured out”), but it is in service of technological development. The author’s admittedly “narrow” portrayal of human history is exclusively focused on the role of science and technology (“compounding scientific discovery and technological progress”). The use of expressive words (“astonishing,” “extraordinarily”) highlights the author’s marvel of technological achievement. The sentence is long and cumulative, depicting a linear progress path culminating in AI. The narrative around “compounding” progress creates a sense of inevitability and inherent momentum in technological development. Finally, the third sentence builds directly on the previous one, underscoring AI’s unparalleled importance and implicitly justifying the immense resources directed towards it. Even though a modal verb (“may”) softens the claim, its superlative nature powerfully portrays technology as the ultimate driver of progress on a simplistic linear path.
@@ 4.1.1 AI-Driven Utopia
This discursive strategy creates a compelling narrative. It first portrays AI as the inevitable culmination of technological progress, and then heralds an era of unparalleled, universally shared prosperity driven by that progress. This dual assertion of historical inevitability and utopian promise functions to legitimize current AI development and deployment paradigms, justifying immense investment and focus. Additionally, this narrative can be used to marginalize critical perspectives or calls for caution as “anti-progress” or unduly pessimistic (cf. “AI doomers”). In doing so, it subtly downplays the intricate requirements of social adaptation, robust governance, and inclusive human agency essential for the responsible development of AI.
4.1.2 Pro-TSD Risk Acknowledgment
While explicit acknowledgment of risk is the most common anti-TSD response to concerns, the data reveal that it was never the final word in any of the analyzed texts. In 13 of the 14 texts (92%), risk acknowledgment was followed by a discursive move aimed at managing, justifying, or resolving the concern in a pro-TSD manner (the single text where this did not occur is andreessen-2023-manifesto because it acknowledges no concerns at all).
This discursive strategy is exemplified by taneja-2022-consequences. This text is part of a book that aims to redefine how technology is developed and invested in, moving towards a more responsible model in which technology must be ethically sound and beneficial. It is part of a narrative ecosystem responding to growing public and regulatory concerns about the negative impacts of technology, offering forward-looking, constructive paths that align technological development with societal good (e.g., pichai-2020-regulate, amodei-2023-scaling, suleyman-2024-containment). Consider the following excerpt:
Throughout history, technology has often been great. But now it must also be good.
Companies that create technology that does good and avoids harmful, unintended consequences will win in the years to come. Investing for financial returns and for positive impact are no longer two different things. They are now the same thing.
Today’s powerful technologies have too much destructive potential […] Now, imagine how Moore’s law and Metcalfe’s law could accelerate the damage done by a biased or rogue artificial intelligence. (Taneja and Maney, 2022, p. 16)
The first paragraph establishes the core message: it is no longer adequate for technology to be merely “great;”; it now needs to be “good.” This paragraph, by distinguishing between greatness and goodness, acknowledges that the previous state of affairs was lacking. The author presents technology in a positive light (“often been great”), but this is immediately followed by a crucial turning point (“But”) emphasized by the juxtaposition of the past (“Throughout history”) and the present (“now”). The use of the modal verb “must” expresses necessity. Both sentences are declarative and straightforward, conveying authority and conviction. The second paragraph claims that companies will be incentivized to “create technology that does good,” not only on moral grounds but also on financial and business grounds. This idea creates a business case for responsible innovation, where responsible companies can still “win,” countering narratives that portray social responsibility as fundamentally in tension with financial success. The alignment between moral responsibility and capitalist incentives is emphasized through the use of rewording (“are no longer two different things,” “They are now the same thing”). This alignment makes the author’s argument palatable to a business and investment audience. Agency is explicitly given to companies (“Companies that create”), and a new criterion for successful technology is established (“technology that does good and avoids harmful, unintended consequences”) using the future tense (“will”) to convey conviction and authority. The third paragraph acknowledges that current technologies have “too much destructive potential” and highlights the potential harm that “a biased or rogue” AI could inflict. This paragraph validates the proposed strategy and portrays the shift from mere greatness to goodness as necessary and urgent. The imperative mood (“Imagine”) directly asks the reader to consider the severity of the risk. While the use of a modal verb (“could”) softens the certainty of the threat, mentioning that Moore’s and Metcalfe’s laws are “accelerating damage” provides concrete, technologically grounded reasons for the urgency.
@@ 4.1.2 Pro-TSD Risk Acknowledgment
The combination of risk acknowledgment with the advancement of a technology-based solution is particularly insidious. The initial acknowledgment validates the reader’s concerns, but the immediate pivot to technological fixes implicitly frames these risks as ultimately problems that can be solved by further technological innovation. The insidious nature of this type of discourse lies in its power of naturalization, a key mechanism of ideological power in Fairclough’s framework. By framing the conversation as ‘risk acknowledged, tech solution proposed,’ the discourse makes the pivot to a technological fix seem commonsensical and natural. When repeated over time, this pattern naturalizes the underlying assumption that the appropriate response to technological harm is always more or better technology, effectively marginalizing non-technological solutions until they seem impractical or irrelevant. The ideology becomes most potent when it is no longer perceived as ideology, but simply as the way things are naturally done. Truly anti-TSD discourse (e.g., BTO) would not only acknowledge risk but also question core TSD tenets and comprehensively explore diverse solutions, including non-tech proposals.
4.2 The Dynamics of TSD Discourse
@@
Figure 3 shows TSD Adherence over time. The chart is contextualized with key milestones in technological capability (e.g., the launch of ChatGPT), foundational ethical and policy framings (e.g., the EU AI Act), and international governance dialogues (e.g., the AI Safety Summit). More details on this broader context are provided in the following section, which serves as a backdrop for the subsequent analysis.
The data reveal distinct patterns of discursive change over time, with most authors demonstrating significant shifts toward more pronounced and imbalanced pro-TSD stances. For example, Pichai moved from a balanced, slightly anti-TSD position in 2020 to a strongly pro-TSD, imbalanced position in 2025. Similarly, Khosla shifted from a highly balanced, neutral stance in 2017 to a strongly pro-TSD, imbalanced stance in 2024. This dynamic can be rapid; within the same year, Altman moved from a moderately pro-TSD text to one that is extremely pro-TSD and imbalanced.
Beyond these individual trajectories, a broader trend of polarization emerges across the corpus, particularly in the period following the launch of ChatGPT. The period from late 2023 onwards exhibits the most extreme TSDA scores in both positive and negative directions. Both the highest pro-TSD scores observed in the corpus (altman-2024-intelligence, andreessen-2023-manifesto, khosla-2024-utopia, pichai-2025-innovation) and the most significantly anti-TSD scores (amodei-2023-scaling, suleyman-2024-containment) are all concentrated in this recent period. Overall, five of the seven authors (71%) showed an increase in their TSDA scores over time, with four of these five (80%) also becoming less balanced. Suleyman is the only author whose TSDA score decreases (and TSDB score increases) after the launch of ChatGPT.
@@
The remainder of the section provides a qualitative analysis of the recent surge of pro-TSD discourse. The broader societal context in which relevant texts were produced is first described. Then, a discussion on potential social determinants contributing to this discursive shift is offered.
4.2.1 From Safety to Action
@@
The period following ChatGPT’s launch in late 2022 has witnessed a dramatic acceleration in AI development and a shift in the dominant discourse among the elite. While the initial generative AI boom was characterized by a robust debate on long-term safety and ethics, this was soon challenged and eclipsed by a powerful counter-narrative emphasizing rapid innovation and geopolitical competition. The new American administration epitomized this shift at the global AI Action Summit in early 2025, where Vice President James David Vance explicitly declared that he was there to discuss opportunity, not safety (34).
@@
As AI capabilities surged after the publication of the Transformer architecture in 2017, the demand for ethical guidelines and regulatory frameworks intensified. Building on a decade of several high-level and voluntary ethics principles from corporations (e.g., Google’s (Pichai, 2018)) and organizations (e.g., OECD’s (1)), the European Union AI Act, first proposed in 2021, signaled a concrete move towards binding regulation (14). The unprecedented user adoption of ChatGPT put AI into the global public consciousness. Throughout 2023, media coverage increasingly fixated on concerns such as algorithmic bias, the dangers of misinformation, and AI-driven job displacement. These anxieties were amplified by seminal figures like Geoffrey Hinton, who resigned from Google to warn of long-term risks, and by the Future of Life Institute’s open letter, urging to “immediately pause the training of AI systems more powerful than GPT-4” in March 2023 (29). In October 2023, then US President Joseph Biden issued a comprehensive Executive Order (36), focusing on AI safety ahead of the first global AI Safety Summit hosted by the United Kingdom in November. The summit culminated in the Bletchley Declaration, in which 28 countries, including the US and China, acknowledged the potential for “serious, even catastrophic, harm” and called for international cooperation (44). The subsequent AI Seoul Summit in May 2024 saw 16 leading AI companies make AI safety commitments; however, it subtly foreshadowed the changing consensus by explicitly adding innovation to the global agenda.
@@
A confluence of powerful technological, commercial, and political forces fueled the discursive shift towards action and innovation. Technologically, the past several years have witnessed an explosion of breakthroughs, including the development of increasingly powerful hardware (e.g., Nvidia’s GPU architecture series), models (e.g., OpenAI’s GPT series; Google’s BERT, Bard, LaMDA, and Gemini series; Anthropic’s Claude series), and practical applications (e.g., DeepMind’s AlphaFold series). Commercially, these technological achievements have attracted unprecedented venture capital, with global AI funding surging past \$100 billion in 2024 (Glaser and Yao, 2025). Additionally, there have been high-profile strategic alliances between companies (e.g., Microsoft’s deepened partnership with OpenAI and Google and Amazon’s investments in Anthropic) to ensure market dominance in a hyper-competitive environment. Politically, competition with China has instilled a sense of urgency for Western innovation, which has been intensified by the release of high-capability open-source models (e.g., Meta’s Llama series, xAI’s Grok series) and the launch of the R1 model by the Chinese company DeepSeek in early 2025. In the United States, Donald Trump’s election marked a pivotal moment. His administration swiftly replaced the previous safety-focused Executive Order with a directive aimed at “removing barriers to American leadership in AI” (35). Moreover, it publicly backed an OpenAI-led investment of \$500 billion in AI (8). This proactive stance garnered international support at the AI Action Summit, hosted by France in February 2025. This summit solidified a shift in global consensus, redirecting AI governance discussions from a predominantly safety-based agenda to one that champions action, opportunity, and accelerated innovation.
4.2.2 The Elite’s Discursive Shift